# Rank distribution in cubic twist families of elliptic curves

### v1

@@ 1.1. Motivation and historical background
Given an elliptic curve $E$ over $\mathbb{Q}$, and a subgroup $H$ of $\operatorname{Aut}_{\bar{\mathbb{Q}}}(E)$, it is natural to consider the twist family of elliptic curves $E_{\zeta}$ indexed by cohomology classes in
$\zeta\in H^{1}(\operatorname{Gal}(\bar{\mathbb{Q}}/\mathbb{Q}),H).$
When $H=\{\pm 1\}$, the associated twists are called quadratic twists and are in bijection with ×/(×$\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$. Goldfeld’s conjecture predicts that the average rank of quadratic twists $E_{m}$ over $\mathbb{Q}$ is $1/2$. In fact, the conjecture predicts that when $m$ is ordered according to its absolute value, exactly half of the quadratic twists have rank 00 and the other half have rank $1$. The rank distribution in quadratic twist families is studied for instance in [BM92, HB94, KMR13, Smi17, KL19, BKOS19]. Set $K:=\mathbb{Q}(\mu_{3})$ and let $\mathcal{O}_{K}$ be its ring of integers. We consider certain families of elliptic curves with rational $3$-isogeny that have complex multiplication by $\mathcal{O}_{K}$. For a sixth power free integer $a\neq 0$, set $E_{a}$ to be the elliptic curve
@@ 1.2. Main results
Let $a$ be an integer and $E_{a}$ be the elliptic curve defined by $y^{2}=x^{3}+a$. Assume that the following conditions are satisfied
@@
Set $\mathcal{Q}_{a}$ to denote the set of primes $\ell\nmid a$ such that the following conditions are satisfied
@@ 1.3. Methodology
The study of Selmer ranks in twist families is related to the study the rank growth and stability of elliptic curves defined over a number field in a $\mathbb{Z}/p\mathbb{Z}$-extension for an odd prime $p$. Let $K$ be a number field containing the $p$-th roots of unity and $E$ be an elliptic curve over $K$ such that the $p$-Selmer group $\operatorname{Sel}_{p}(E/K)=0$. Let $L/K$ be a cyclic $p$-extension. Then, it is shown that under certain conditions on $E$ and $L$, the Selmer group $\operatorname{Sel}_{p}(E/L)=0$ as well. This strategy has been adapted from the technique of Mazur–Rubin [MR07, MRS07], Brau [Bra14] and Česnavičius [Č17]. In keeping with these works, we study stability questions in prime cyclic extensions for any prime number $p\geq 3$, cf. Theorem 3.12 and Theorem 3.15 for further details.
@@ 1.4. Related work
In recent years, there has been an increasing interest in rank distribution questions for families of elliptic curves from the perspective of arithmetic statistics. These techniques involve the parameterization of Selmer groups of elliptic curves and abelian varieties by integral forms in many variables. The distribution properties of such integral forms can be studied via techniques from analytic number theory and the geometry of numbers, and has gained momentum in works of Bhargava–Shankar [BS15a, BS15b, BS13], Bhargava–Elkies–Shnidman [BES20], Alpöge–Bhargava–Shnidman [ABS22], Shnidman–Weiss [SW21, SW23] etc. Our techniques on the other hand rely of local and global arguments in Galois cohomology (in keeping with those used by Mazur–Rubin, Brau, Česnavičius), and one of the consequences is that we obtain very explicit descriptions for the twist families for which the Selmer group group vanishes. It is conceivable that our results should generalize to the context of abelian varieties with complex multiplication. For instance, Jacobians of twist families of bicyclic trigonal curves $y^{3}=f(x^{2})$, cf. [SW21] for details.
@@ 1.5. Organization
Including the introduction, the article consists of four sections. In section 2, we set basic notation and recall definitions of Selmer groups and their properties. In section 3, we study the stability of $p$-Selmer groups in $p$-cyclic extensions of number fields, leveraging the technique of Mazur–Rubin, Brau and Česnavičius. We prove our main results for cubic twist families in section 4. These results are backed up by explicit numerical computations in subsections 4.3 and 4.4.
@@ 2. Notation and preliminaries
Let $E$ be an elliptic curve over $K$ and $n$ be a positive integer. Given a prime $v\in\Sigma_{K}$ at which $E$ has good reduction, set $\widetilde{E}$ to denote the reduction of $E$ at $v$. Take $\widetilde{E}(k_{v})$ to denote the group of $k_{v}$-points on $\widetilde{E}$. Set $E[n]$ to denote the $n$-torsion subgroup if $E(\bar{K})$. For $\mathcal{F}\in\{K,K_{v}\}$, we have the exact sequence of $\operatorname{G}_{\mathcal{F}}$-modules
@@ 3.1. Local contributions to the growth of Selmer groups
We take $L/K$ to be a finite Galois extension with Galois group $G=\operatorname{Gal}(L/K)$ isomorphic to $\mathbb{Z}/p\mathbb{Z}$. Let $\sigma$ be a generator of $G$. Set $S$ to be the set of primes $v\in\Sigma_{K}$ such that either one of the following conditions are satisfied
@@
Denote by $E_{1}(K_{v})$ (resp. $E_{1}(L_{w})$) denote the kernel of the reduction map $E(K_{v})\rightarrow\widetilde{E}(k_{v})$ (resp. $E(L_{w})\rightarrow\widetilde{E}(k_{w})$). Since $\widetilde{E}(k_{v})[p]=0$ and $L/K$ is a $p$-extension, it follows that $\widetilde{E}(k_{w})[p]=0$ (cf. Lemma 3.1). Therefore, we find that
@@
Since $p\nmid\widetilde{E}(k_{w})$, it follows from the restriction-corestriction sequence that
@@
Assume without loss of generality that $v$ is inert or ramified in $L$. Let $w^{\prime}$ be the prime of $L^{\prime}$ that lies above $v^{\prime}$. Setting $L^{\prime}:=K^{\prime}\cdot L$, let $w$ be a prime of $L$ that lies below $w^{\prime}$, as depicted
@@ 3.2. Rank stability in prime cyclic extensions
Let $K$ be a number field and $E$ be an elliptic curve over $K$. Let $p$ be an odd prime number and assume that $\operatorname{Sel}_{p}(E/K)=0$. Let $L/K$ be a degree $p$ Galois extension and $S$ be the set of primes $v$ of $K$ such that either one of the following hold
@@ 4.2. Main results
$\operatorname{det}\rho_{E,\ell}(\operatorname{Frob}_{\ell})=\ell\text{ and }\operatorname{trace}\rho_{E,3}(\operatorname{Frob}_{\ell})=\ell+1-\#\widetilde{E}(\mathbb{F}_{\ell}).$

### v2

@@ Abstract.
Key words and phrases:
Elliptic curves, cubic twist families, rank distribution, Selmer groups and stability, explicit computation
@@ 1.1. Motivation and historical background
Given an elliptic curve $E$ over $\mathbb{Q}$, and a subgroup $H$ of $\operatorname{Aut}_{\bar{\mathbb{Q}}}(E)$, it is natural to consider the twist family of elliptic curves $E_{\zeta}$ indexed by cohomology classes $\zeta\in H^{1}(\operatorname{Gal}(\bar{\mathbb{Q}}/\mathbb{Q}),H)$. When $H=\{\pm 1\}$, the associated twists are called quadratic twists and are in bijection with ×/(×$\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$. Goldfeld’s conjecture predicts that the average rank of quadratic twists $E_{m}$ over $\mathbb{Q}$ is $1/2$. In fact, the conjecture predicts that when $m$ is ordered according to its absolute value, exactly half of the quadratic twists have rank 00 and the other half have rank $1$. The rank distribution in quadratic twist families is studied for instance in [BM92, HB94, KMR13, Smi17, KL19, BKOS19]. Set $K:=\mathbb{Q}(\mu_{3})$ and let $\mathcal{O}_{K}$ be its ring of integers. We consider certain families of elliptic curves with rational $3$-isogeny that have complex multiplication by $\mathcal{O}_{K}$. For a sixth power free integer $a\neq 0$, set $E_{a}$ to be the elliptic curve
@@ 1.2. Main results
Let $a$ be an integer and $E_{a}$ be the elliptic curve defined by $y^{2}=x^{3}+a$. Assume that the following conditions are satisfied:
@@
Set $\mathcal{Q}_{a}$ to denote the set of primes $\ell\nmid a$ such that the following conditions are satisfied:
@@ 1.3. Methodology
The study of Selmer ranks in twist families is related to the study the rank growth and stability of elliptic curves defined over a number field in a $\mathbb{Z}/p\mathbb{Z}$-extension for an odd prime $p$. Let $K$ be a number field containing the $p$-th roots of unity and $E$ be an elliptic curve over $K$ such that the $p$-Selmer group $\operatorname{Sel}_{p}(E/K)=0$. Let $L/K$ be a cyclic $p$-extension. Then, it is shown that under certain conditions on $E$ and $L$, the Selmer group $\operatorname{Sel}_{p}(E/L)=0$ as well. This strategy has been adapted from the technique of Mazur–Rubin [MR07, MRS07], Brau [Bra14]. In keeping with these works, we study stability questions in prime cyclic extensions for any prime number $p\geq 3$, cf. Theorem 3.12 and Theorem 3.15 for further details.
@@ 1.4. Related work
In recent years, there has been an increasing interest in rank distribution questions for families of elliptic curves from the perspective of arithmetic statistics. These techniques involve the parameterization of Selmer groups of elliptic curves and abelian varieties by integral forms in many variables. The distribution properties of such integral forms can be studied via techniques from analytic number theory and the geometry of numbers, and has gained momentum in works of Bhargava–Shankar [BS15a, BS15b, BS13], Bhargava–Elkies–Shnidman [BES20], Alpöge–Bhargava–Shnidman [ABS22], Shnidman–Weiss [SW21, SW23] etc. Our techniques on the other hand rely on local and global arguments in Galois cohomology (in keeping with those used by Mazur–Rubin, Brau), and one of the consequences is that we obtain very explicit descriptions for the twist families for which the Selmer group vanishes. Related questions have been studied by Satgé [Sat86], Iskra [Isk96], Goto [Got01] and Bandini [Ban04]. It is conceivable that our results should generalize to the context of abelian varieties with complex multiplication. For instance, Jacobians of twist families of bicyclic trigonal curves $y^{3}=f(x^{2})$, cf. [SW21] for details.
@@ 1.5. Organization
Including the introduction, the article consists of four sections. In section 2, we set basic notation and recall definitions of Selmer groups and their properties. In section 3, we study the stability of $p$-Selmer groups in $p$-cyclic extensions of number fields, leveraging the techniques of Mazur–Rubin, Brau and Česnavičius. We prove our main results for cubic twist families in section 4. These results are backed up by explicit numerical computations in subsections 4.3 and 4.4.
1.6. Acknowledgement
Thanks to Antonio Lei for answering some of our questions, and for helpful comments. We would like to thank the referee for the careful review of the manuscript.
Data availability
No data was analyzed in proving the results in the article.
Conflict of Interest
There is no conflict of interest that the authors wish to report.
@@ 2. Notation and preliminaries
Let $E$ be an elliptic curve over $K$ and $n$ be a positive integer. Given a prime $v\in\Sigma_{K}$ at which $E$ has good reduction, set $\widetilde{E}$ to denote the reduction of $E$ at $v$. Take $\widetilde{E}(k_{v})$ to denote the group of $k_{v}$-points on $\widetilde{E}$. Set $E[n]$ to denote the $n$-torsion subgroup of $E(\bar{K})$. For $\mathcal{F}\in\{K,K_{v}\}$, we have the exact sequence of $\operatorname{G}_{\mathcal{F}}$-modules
@@ 3.1. Local contributions to the growth of Selmer groups
We take $L/K$ to be a finite Galois extension with Galois group $G=\operatorname{Gal}(L/K)$ isomorphic to $\mathbb{Z}/p\mathbb{Z}$. Let $\sigma$ be a generator of $G$. Set $S$ to be the set of primes $v\in\Sigma_{K}$ such that any one of the following conditions are satisfied:
@@
Denote by $E_{1}(K_{v})$ (resp. $E_{1}(L_{w})$) the kernel of the reduction map $E(K_{v})\rightarrow\widetilde{E}(k_{v})$ (resp. $E(L_{w})\rightarrow\widetilde{E}(k_{w})$). Since $\widetilde{E}(k_{v})[p]=0$ and $L/K$ is a $p$-extension, it follows that $\widetilde{E}(k_{w})[p]=0$ (cf. Lemma 3.1). Therefore, we find that
@@
Since $p\nmid\#\widetilde{E}(k_{w})$, it follows from the restriction-corestriction sequence that
@@
Assume without loss of generality that $v$ is inert or ramified in $L$. Setting $L^{\prime}:=K^{\prime}\cdot L$, let $w^{\prime}$ be the prime of $L^{\prime}$ that lies above $v^{\prime}$. Let $w$ be a prime of $L$ that lies below $w^{\prime}$, as depicted
@@ 3.2. Rank stability in prime cyclic extensions
Let $K$ be a number field and $E$ be an elliptic curve over $K$. Let $p$ be an odd prime number and assume that $\operatorname{Sel}_{p}(E/K)=0$. Let $L/K$ be a degree $p$ Galois extension and $S$ be the set of primes $v$ of $K$ such that any one of the following hold:
@@ 4.2. Main results
$\operatorname{det}\rho_{E,3}(\operatorname{Frob}_{\ell})=\ell\text{ and }\operatorname{trace}\rho_{E,3}(\operatorname{Frob}_{\ell})=\ell+1-\#\widetilde{E}(\mathbb{F}_{\ell}).$